Week of 201130

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local:
  • remote: Christoph (CMS), Dave (FNAL), Maarten (ALICE), Miro (Chair, DB), Onno (NL-T1), Andrew (TRIUMF), Xavier (KIT), Alberto (Monitoring), David (IN2P3), Vincent (Security), Vladimir (LHCb), Elena (CNAF), Darren (RAL) , Xin (BNL), Ivan (ATLAS)

Experiments round table:

  • CMS reports ( raw view) -
    • Recend update of HammerCloud
      • Production CMS instance got updated on Thursday without prior announcement
      • Severe problems since then
      • Request a roll back and proper validation in test environment
    • Very good CPU utilization (beyond 300k cores in the CMS Global Pool)

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • PIC dCache namespace migration today without downtime
    • Activity:
      • MC and WG productions; User jobs.

Sites / Services round table:

  • ASGC: NC
  • BNL: a reminder of the upcoming downtime (Dec. 1st) at the Tier1, for network upgrade.
  • EGI: NC
  • IN2P3:
    • LHCb: name space reordering done on last Tuesday Nov. 24th
    • site in downtime on Monday 7th end of day and Tuesday 8th the whole day. Electric power supply maintenance. CEs and SEs in downtime.
  • JINR: NC
  • KIT: NTR
  • NDGF: NC
  • NL-T1: Tomorrow Surfsara's tape system will be in maintenance. https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=29817
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: NTR

  • CERN computing services: NC
  • CERN storage services: NC
  • CERN databases: NTR
  • Monitoring: NTR
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR


Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2020-11-30 - ChristophWissing
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback